# Routine haematological markers can predict and discriminate health status and biological age even from noisy sources

### v10

@@ 3.1 Adaptive Index Behaviour Over Time
In 2, we tested the behaviour of the NIS index over time by simulating patients evolving in different directions with outliers. We show that the index can be used to learn setpoints or baselines of patients over time based on the behaviour of the index flagging for changes even within the normal values of the population.
@@
The results show significantly different behaviour over time for male and female cohorts. For subjects over the age of 30, it is evident that the male population presents an upward trajectory with respect to age for the mean immune index, as well as an increase in the 50th percentile and variance. The immune index decreases slightly for males between 20 and 25 year old. However, the trend in women slows down (compared to men) between ages 40 and 50 and reverses between ages 50 and 60. It remains to investigate whether some of the analytes involved are confounded with contradictory trends from other physiological differences, such as menopause, which occurs exactly between the ages when the slowdown and reversal of the age index trend for women is detected. Pregnancy and menopause are known to produce a more progressive and higher increase in red blood cell counts, hemocrit levels, and increased mean cell volume (MCV), as well as hemoglobin concentrations [41, 19, 24, 44, 13].
@@ 3.2 Discounting Confounding Factors from Age
where $\alpha^{*}$ is the closest index in $IAC$ to $\alpha$. If there are two scores at the same distance, we choose the leftmost option. The resulting function is shown in Fig.  4. The curve shows some minor bumps resulting from the discretisation used. We maintain that this distortion is not significant.
As a tool that returns an estimated age with respect to a index, we can analyse its predictive power. Using the NHANES data and focusing on males aged 20 to 90, we can measure the error of using the index to predict a person’s age. As shown in Figure  5, prediction error rates are much lower in healthy NHANES data, compared to unhealthy groups, for both sexes, and therefore can be used to estimate the chronological age of healthy patients. Prediction error rates for the immune index age were overestimated by approximately 7 years, for the healthy groups of both sexes, allowing a correction (mean error rate of 7.33 $\pm$ 1.21 for healthy females, and 7.62 $\pm$ 1.17 for healthy males). Furthermore, there is a statistically significant difference between the healthy and non-healthy groups (regardless of sex) in the error rates of age prediction, determined by two-tailed, two-sample t-tests. Pearson’s correlation statistics on the age error rates predicted by the health status of the red blood cell count as discussed before, were an R square value of 0.293 and 0.168, for men (P<< 0.0001) and women (P<<0.013) respectively. Age prediction error rates were high and not significant when self-reported health measures were used to separate health status groups, but were significant when taking abnormal CBC values only.
@@ 3.4 UK Biobank Subset Validation
We first filtered the Rest (‘healthy’ control) cohort by removing all individuals whose third-time point CBC measurements deviated from established normal physiological ranges. This reduced the noise and the cohort size from $n=2312$ to $n=1741$. For disease groups (Cancer, CVD, Infection), we kept only those individuals who exhibited at least four CBC values that fall outside the normal range. This filtering reduced the initial cohort sizes (Cancer: 666, CVD: 632, Infection: 396) to 438, 417, and 367 respectively. Pairwise comparisons of composite NIS scores between these filtered groups are summarised in Table 13, and box plots of scores are presented in Fig. 7. The Shapiro–Wilk tests on the residuals yielded p-values << 0.05 for all groups, thus we used the Kruskal–Wallis test followed by Bonferroni-corrected Mann–Whitney U tests for pairwise comparisons, as shown in Table 13.
@@ 4 Conclusions
We demonstrated the capabilities and digital applications of the scores for triaging purposes against two public databases, showing how the scores separate healthy from unhealthy cohorts, both as self-reported, with significant noise, and as determined by typical normal reference values. The approach was shown to be informative for quick and entry-level patient health sorting, screening and triaging.
@@
From an implementation standpoint, these indices could be embedded into electronic health record (EHR) systems as automated warning or longitudinal monitoring modules, providing clinicians with real-time quantitative feedback on patient trajectories. Similar to widely used risk assessment scores, the proposed metric condenses deviations in blood test parameters into a single, symbolically interpretable, and directly comparable value for applications such as triage, where the alternative would be a first-come, first-served or random-order approach.

### v12

@@ 3.1 Adaptive Index Behaviour Over Time
In 3, we tested the behaviour of the NIS index over time by simulating patients evolving in different directions with outliers. We show that the index can be used to learn setpoints or baselines of patients over time based on the behaviour of the index flagging for changes even within the normal values of the population.
@@
The results show significantly different behaviour over time for male and female cohorts. For subjects over the age of 30, it is evident that the male population presents an upward trajectory with respect to age for the mean immune index, as well as an increase in the 50th percentile and variance. The immune index decreases slightly for males between 20 and 25 year old. However, the trend in women slows down (compared to men) between ages 40 and 50 and reverses between ages 50 and 60.
This observed inflection in females is reported descriptively rather than causally interpreted. While menopause represents a temporally coincident physiological transition that may influence haematological parameters, the present dataset does not include menopause status, menopause-specific hormonal measurements, or neuro-immune-endocrine axes required for mechanistic validation.
Therefore, the association should be considered as suggestive systemic explanations or prospective study-generating hypotheses. Pregnancy and menopause have been associated with alterations in red blood cell counts, haematocrit levels, mean cell volume (MCV), and haemoglobin concentrations [48, 26, 31, 51, 20], suggesting plausible mechanistic pathways.
However, given the multiscale physiological complexity underlying these processes, dedicated stratified longitudinal analyses incorporating reproductive and hormonal status would be necessary to formally validate this interpretation.
@@ 3.2 Discounting Confounding Factors from Age
where $\alpha^{*}$ is the closest index in $IAC$ to $\alpha$. If there are two scores at the same distance, we choose the leftmost option. The resulting function is shown in Fig.  5. The curve shows some minor bumps resulting from the discretisation used. We maintain that this distortion is not significant.
As a tool that returns an estimated age with respect to a index, we can analyse its predictive power. Using the NHANES data and focusing on males aged 20 to 90, we can measure the error of using the index to predict a person’s age. As shown in Figure  6, prediction error rates are much lower in healthy NHANES data, compared to unhealthy groups, for both sexes, and therefore can be used to estimate the chronological age of healthy patients. Prediction error rates for the immune index age were overestimated by approximately 7 years, for the healthy groups of both sexes, allowing a correction (mean error rate of 7.33 $\pm$ 1.21 for healthy females, and 7.62 $\pm$ 1.17 for healthy males). Furthermore, there is a statistically significant difference between the healthy and non-healthy groups (regardless of sex) in the error rates of age prediction, determined by two-tailed, two-sample t-tests. Pearson’s correlation statistics on the age error rates predicted by the health status of the red blood cell count as discussed before, were an R square value of 0.293 and 0.168, for men (P<< 0.0001) and women (P<<0.013) respectively. Age prediction error rates were high and not significant when self-reported health measures were used to separate health status groups, but were significant when taking abnormal CBC values only.
We emphasise that the immune age metric is not intended to function as a high-precision ’chronological’ ageing clock comparable to existing methods, such as epigenetic or proteomic biological age estimators (e.g., methylation-based clocks).
Rather, its purpose is to detect divergence between routinely accessible physiological immune patterns and expected age-related norms derived from CBC blood counts.
The observed mean error largely reflects systematic offset rather than random dispersion. Post-correction bias is reduced and variance narrows substantially, particularly within healthy cohorts. Importantly, prediction error is lower in physiologically stable (healthy) groups, reinforcing its role in deviation detection rather than exact age estimation.
Accordingly, the index is positioned as a population-level screening and stratification tool, not a diagnostic clock, designed to support accessible, affordable, and equitable health assessment with translation towards personalised patient care.
@@ 3.3 Machine Learning Guiding index
The interpretability of the hybrid neurosymbolic approach emerges from the fact that the neural networks function on a single scalar immune index (NIS) rather than on high-dimensional raw feature vectors.
Feature importance rankings derived from Random Forest models (mean Gini decrease; Table 9) were used to identify the most influential CBC analytes contributing to discrimination, providing an interpretable predictive ranking of variables.
The neural networks were then applied either to the full analyte set or directly to the scalar NIS score (Tables 10–12). Thus, when operating on the NIS, the neural component refines decision boundaries within a symbolically defined biological parameter space, rather than discovering latent representations de novo.
This preserves explainability of the contributing variables while allowing nonlinear pattern detection through computational intelligence. Future extensions of this neurosymbolic AI framework could incorporate architecture-dependent feature pruning, knowledge destination with hybrid combination of architectures, sparsity-constrained optimisation, or regularisation-driven layer compression to further align neural models’ predictive performance with biologically meaningful features identified through neurosymbolic ranking.
@@ 3.4 UK Biobank Subset Validation
Furthermore, the availability of repeated CBC measurements in the UK Biobank subset allows partial real-world assessment of temporal robustness related to the adaptive baseline introduced in Fig. 2. While the adaptive learning itself was illustrated through simulation, the averaging of three longitudinal timepoints prior to scoring demonstrates that the immune index retains discriminatory power under real-world temporal variability. The primary objective of this study is to demonstrate that the immune index discriminates health status and approximates biological age using routine CBC measurements, with multi-timepoint averaging in the UK Biobank serving as a robustness validation rather than the central methodological requirement. Thus, the index remains stable and predictive when applied to multi-timepoint clinical data.
@@
We first filtered the Rest (‘healthy’ control) cohort by removing all individuals whose third-time point CBC measurements deviated from established normal physiological ranges. This reduced the noise and the cohort size from $n=2312$ to $n=1741$.
For disease groups (Cancer, CVD, Infection), we kept only those individuals who exhibited at least four CBC values that fall outside the normal range. This filtering reduced the initial cohort sizes (Cancer: 666, CVD: 632, Infection: 396) to 438, 417, and 367 respectively.
Pairwise comparisons of composite NIS scores between these filtered groups are summarised in Table 13, and box plots of scores are presented in Fig. 8.
The Shapiro–Wilk tests on the residuals yielded p-values << 0.05 for all groups, thus we used the Kruskal–Wallis test followed by Bonferroni-corrected Mann–Whitney U tests for pairwise comparisons, as shown in Table 13.
@@
As discussed, future work should also incorporate menopause status and hormonal descriptions available in UK Biobank to formally test whether the observed immune index inflection in females reflects endocrine transition rather than sampling heterogeneity. Such stratification would enable mechanistic refinement of the sex-specific immune ageing curve.
Lastly, while the NHANES cohort includes conditions with pronounced haematological disruption (e.g., leukemia, HIV), the UK Biobank validation extends the analysis to chronic cardiovascular disease and infection cohorts (Table 13). These represent systemic inflammatory and metabolic conditions rather than primary marrow-disruptive pathologies. The preservation of discriminatory performance in these cohorts suggests that the immune index captures broader physiological dysregulation of the blood–immune–inflammatory axis beyond overt haematological malignancy. Future studies may further evaluate the applicability of the immune index in developmental, neuropsychiatric and other systemic inflammatory and metabolic disorders, where immune perturbations are increasingly recognized as contributing factors.
@@ 4 Conclusions
We demonstrated the capabilities and digital applications of the scores for triaging purposes against two public databases, showing how the scores separate healthy from unhealthy cohorts, both as self-reported, with significant noise, and as determined by typical normal reference values.
The approach was shown to be informative for quick and entry-level patient health sorting, screening and triaging.
@@
From an implementation standpoint, these indices could be embedded into electronic health record (EHR) systems as automated warning or longitudinal monitoring modules, providing clinicians with real-time quantitative feedback on patient trajectories.
Similar to widely used risk assessment scores, the proposed metric condenses deviations in blood test parameters into a single, symbolically interpretable, and directly comparable value for applications such as triage, where the alternative would be a first-come, first-served or random-order approach.
@@
Declaration of Competing Interest
The authors declare the following competing financial interests that may be considered as potential competing interests: the authors are affiliated with Oxford Immune Algorithmics, which may commercially develop applications related to the index described in this work under its AI platform for precision healthcare called Algocyte and which they often refer to as Algocyte index.
@@ 4.5 Adaptive Reference Values
Supplementary Methods: Additional Clock Estimators
[…]
Further, we compared our immune score to real-world inflammatory immune ratios used in clinical context, namely: 1) NLR (Neutrophil-to-Lymphocyte Ratio) == Neutrophils / Lymphocytes, 2) PLR (Platelet-to-Lymphocyte Ratio) == Platelets / Lymphocytes, 3) SII (Systemic Immune-Inflammation Index) == Platelets $\times$ Neutrophils / Lymphocytes, and 4) SIRI (Systemic Inflammation Response Index) == Neutrophils $\times$ Monocytes / Lymphocytes. These are low-dimensional ratio-based markers and thus, reduce the immune complexity into 2–3 interacting cell types and are predicted to underperform compared to the multivariate models.
Figure S4: CBC Age Acceleration (DeltaAge) stratified by health and sex. CBC_AgeClock_DeltaAge == Predicted CBC-age - chronological age, where Predicted CBC-age is generated by a ridge-regression CBC age clock ($\alpha\approx 56.23$) trained on healthy subjects using CBC-only features (WBC, RBC, MCV, MCH, MCHC). Unhealthy participants show significantly lower DeltaAge than healthy overall and within each sex. Among unhealthy individuals, males exhibit more negative DeltaAge than females.
Figure S4 shows the delta age from the Ridge Regression-based CBC clock simulation, where delta age is defined as Predicted CBC-age - chronological age, wherein the Predicted CBC-age comes from a ridge-regression CBC age clock ($\alpha\approx 56.23$) trained only on healthy subjects using the shared CBC features (i.e., WBC, RBC, MCV, MCH, MCHC). Positive DeltaAge means the CBC profile appears older than expected for that person’s age (age acceleration), while negative DeltaAge means it appears younger. In the pooled comparison, unhealthy individuals show markedly lower DeltaAge than healthy (Healthy mean 0.00 vs Unhealthy mean 9.79-9.79; Cohen’s $d=0.556$; Welch $p=2.16\times 10^{-56}$). Stratifying by sex clarifies the pattern: within females, Healthy vs Unhealthy remains significant (mean 0.31 vs 7.58-7.58; $d=0.443$; $p=1.17\times 10^{-27}$), and within males the separation is even larger (mean 0.30-0.30 vs 15.70-15.70; $d=0.890$; $p=1.45\times 10^{-38}$). Within the unhealthy group, males are more negative than females (mean 15.70-15.70 vs 7.58-7.58; d=0.439d=-0.439; $p=4.76\times 10^{-11}$), while in healthy subject’s male–female differences are small/borderline ($p\approx 0.055$).
Figure S5: Multivariate CBC deviation from healthy reference stratified by health and sex. Mahalanobis distance was computed in standardized CBC 5-feature space relative to the healthy training distribution. Unhealthy individuals exhibit increased multivariate hematologic deviation compared to healthy subjects, with minimal sex-dependent effects within strata.
Mahalanobis distance (Fig. S5) shows clear separation between healthy and unhealthy individuals as a disease deviation hematologic manifold, with unhealthy subjects exhibiting substantially greater multivariate deviation from the normative hematologic centroid. Across the full cohort, the effect size for Healthy vs Unhealthy is moderate-to-large (Cohen’s $d\approx 0.68$; $p\ll 10^{-30}$), and the separation remains significant within both females ($d\approx 0.77$) and males ($d\approx 0.60$). In contrast, sex differences are negligible within both healthy and unhealthy strata ($p>0.1$ in both cases). Thus, Mahalanobis distance, as a manifold learning algorithm, primarily captures health-related deviation rather than sex-driven differences (i.e., health-state discrimination), whereas the above defined delta age may be more insightful for capturing biological ageing dynamics.
Figure S6: Z-distance from healthy CBC reference stratified by health and sex. ZDistance quantifies Euclidean distance in standardized CBC feature space (WBC, RBC, MCV, MCH, MCHC) from the healthy reference centroid, thereby measuring global hematologic deviation from normative structure. Unhealthy individuals show substantially greater multivariate deviation than healthy subjects (mean 3.005 vs 2.024; median 2.487 vs 1.799), corresponding to a large effect size (Cohen’s $d\approx 0.76$) and highly significant separation (Welch $p\approx 2.1\times 10^{-32}$). This discrimination remains robust within both males and females, while sex differences themselves are comparatively small, indicating that health status and not sex is the dominant driver of variation.
Figure S7: Systemic Immune-Inflammation Index (SII; platelets $\times$ neutrophils / lymphocytes) shows only a very small sex effect (mean 490.2 vs 540.9; Cohen’s $d\approx 0.14$) despite extreme statistical significance from large sample size ($p\approx 4.8\times 10^{-90}$), indicating weak biological separation compared to multivariate measures like Z-distance or Mahalanobis.
Figure S8: Systemic Inflammation Response Index (SIRI; neutrophils $\times$ monocytes / lymphocytes) demonstrates an even smaller sex effect (mean 1.17 vs 1.09; Cohen’s $d\approx 0.08$) despite strong statistical significance from large sample size ($p\approx 1.4\times 10^{-29}$), indicating biologically minimal separation and substantially weaker discriminative value than multivariate measures such as Z-distance or Mahalanobis.
Figure S9: Platelet-to-Lymphocyte Ratio (PLR; platelets / lymphocytes) shows only a small sex effect (Cohen’s $d\approx 0.095$) despite strong statistical significance driven by large sample size, indicating minimal biological separation between males and females. Its discriminative performance is modest—slightly stronger than SIRI and similar to SII—but substantially weaker than multivariate measures such as Mahalanobis distance or Z-distance, which capture coordinated hematologic structure rather than single-ratio shifts.
Figure S10: Neutrophil-to-Lymphocyte Ratio (NLR; neutrophils / lymphocytes) is the weakest of the evaluated indices, with an essentially negligible effect size (Cohen’s $d\approx 0.016$), indicating biologically trivial male–female separation despite statistical significance driven by large sample size. It performs substantially worse than PLR, SII, and SIRI, and far below multivariate approaches such as Mahalanobis distance or Z-distance.
Conclusion
[…]
Further, our immune score outperforms these traditional ratio-based immune markers because it captures global hematologic structure rather than a single inflammatory axis (Figs. S7S10). SII, SIRI, PLR, and NLR are low-dimensional, highly correlated composites derived from two or three cell types, limiting their capacity to detect coordinated immune–hematologic perturbations. By integrating multiple orthogonal CBC features and modeling their joint covariance structure, our score identifies system-level dysregulation rather than isolated ratio shifts, resulting in substantially larger, biologically meaningful effect sizes and clearer discrimination of health status.